A man, stride is cm long and his wife stride is 15 cm shorter. He covers 30m in 10 steps less than his wife. frame an equation in x and solve it
1
Expert's answer
2019-06-13T09:34:28-0400

Let the man's stride be x cm.

Then the wife's stride is (x - 15) cm.


The distance that they cover is 30 m = 3000 cm.

The number of steps taken by man in 30 m is 3000/x.

The number of steps taken by wife in 30 m is 3000/(x - 15).


As the man covers 30 m in 10 steps less than his wife, we get the equation:


3000/(x - 15) - 10 = 3000/x


3000x - 10x(x-15) = 3000(x -15)


3000(x - x + 15) - 10(x2 - 15x) = 0


3000 * 15 = 10(x2 - 15x)


x2 - 15x - 4500= 0


x2 - 75x + 60x - 4500 = 0


x(x - 75) + 60(x - 75) = 0


(x - 75)(x + 60) = 0


x = 75 cm


x - 15 = 60 cm


Thus, the man's stride is 75 cm and the wife's stride is 60 cm.